The burning of 5.08g of C6H6 releases enough heat to raise the temperature of 50g of water from 10.1℃ to 19.6℃. Calculate ΔH/mole.
1
Expert's answer
2020-05-19T08:47:08-0400

CH2O = 4.186 J ∙ mol-1 ∙ K-1

Q = 4.186 ∙ 50 ∙ (19.6 - 10.1) = 1988.35 J

n(C6H6) = 5.08 / 78 = 0.065 mol

ΔH = 1988.35 / 0.065 = 30590 (J / mol) = 30.59 (kJ / mol)

Answer: ΔH = 30.59 (kJ / mol).